If the minimum point of the short run ATC curve for all firms (existing and potential) is also the minimum point of the long run average cost curve (LRAC), calculate the long-run equilibrium price, market quantity, and firm quantity. What is the long-run equilibrium number of firms in the industry?


1
Expert's answer
2021-11-14T17:32:41-0500

There are 100 identical firms ,

the supply curve of the market is given by , Qs = nQ= 100"\times" P = 100P

3) At equilibrium Qs= Qd

100P = 1000 -100P

200P = 1000

P = 5

put this value in either demand or supply function

Q= 1000 - 5"\times" 100

= 1000 - 500

= 500

Equilibrium quantity of the market is 500

Firm Quantity = Market quantity "\div" 100

"=\frac{ 500}{100}"

= 5